Scientific notation for 2002 sales of $58,200,000,000 and 2006 sales of
$90,800,000,000 and percentage of growth?
thank you, Helen

Answer by checkley77(12569)

58,200,000,000=5.82*10^10
90,800,000,000=9.08*10^10
9.08*10^10/5.82*10^10
9.08/5.82=1.56 or 156% increase.